OpenStudy (anonymous):

solve for x: (e^4)^x*e^x^2=e^12

OpenStudy (anonymous):

\[(e ^{4})^{x}\times e ^{x ^{2}}=e ^{12}\]

OpenStudy (anonymous):

did i make it clear now ?

OpenStudy (anonymous):

you have e^(4x+x^2)=e^12 to be equal should x^2+4x=12 solve it and it is

OpenStudy (anonymous):

\[e ^{4x+x ^{2}}=e ^{12}\] by exponent rules later just take logarithm of bouth sides and solve equation for x

OpenStudy (anonymous):

oh thanks. i got it. can you help me with this aswel ? \[e ^{4x-9}=\frac{ 1 }{ e }\]

OpenStudy (anonymous):

here u have 4x-9=-1

OpenStudy (anonymous):

\[1/e=e ^{-1}\] the rest is the same like in previous one

OpenStudy (anonymous):

bec0use 1/e=e^-1

OpenStudy (anonymous):

oh thank you guys :) appreciate it a lot